Two dice are rolled. What is the probability of scoring a double given that the sum on the two dice is less than 6?


1
Expert's answer
2021-09-13T00:05:11-0400

Sum table



Total number of variants N = 36

A "double" means that two dice have the same number.

Positive combinations when the sum on the two double dice is less than 6: (1,1) and (2,2).

n=2

The probability of scoring a double given that the sum on the two dice is less than 6:

"P = \frac{n}{N}= \frac{2}{36}= \frac{1}{18}"

Answer: "\frac{1}{18}"